# Pipeline Gradient-based Model Training on Analog In-memory Accelerators

### v1

@@ Appendix E Proof of Theorem 2: Analog SGD with asynchronous pipeline
$\ -\left(\frac{1}{4\alpha}-\alpha(1+\frac{1}{u})(1+W_{\max}/\tau)^{2}C_{f}^{2}M^{2}\right)\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]$
$\ +\alpha^{3}(1+\frac{1}{u})(1+W_{\max}/\tau)^{2}C_{f}^{2}M^{3}\sigma^{2}.$
Taking expectation, averaging (93) over $k$ from 00 to $K-1$, and choosing the learning rate $\alpha\leq\frac{1}{2\sqrt{1+1/u}(1+W_{\max}/\tau)C_{f}M}$ yield
@@
$\leq$ $\ \frac{2({\mathbb{V}}_{0}-{\mathbb{E}}[{\mathbb{V}}_{k+1}])}{\alpha K(1-(1+u)W_{\max}^{2}/\tau^{2})}+\frac{2\alpha L_{f}\sigma^{2}}{1-(1+u)W_{\max}^{2}/\tau^{2}}$
$\ +\sigma^{2}\frac{1}{K}\sum_{k=0}^{K-1}\frac{(1+u)\|W_{k}\|^{2}_{\infty}/\tau}{1-(1+u)\|W_{k}\|^{2}_{\infty}/\tau^{2}}+\frac{\alpha^{2}(1+1/u)(1+W_{\max}/\tau)^{4}C_{f}^{2}M^{3}\sigma^{2}}{(1-(1+u)W_{\max}^{2}/\tau^{2})}.$
@@
$\leq$ $\ 4\sqrt{\frac{(f(W_{0})-f^{*})\sigma^{2}L_{f}}{K}}\frac{1}{1-(1+u)W_{\max}^{2}/\tau^{2}}+\sigma^{2}S_{K}^{\prime}+O\left(\frac{1+1/u}{K}\right).$
@@
Appendix F Convergenc of digital asynchronous pipeline SGD
Theorem 3 (Convergence rate, asynchronous pipeline).
@@ Appendix F Convergenc of digital asynchronous pipeline SGD
[…]
@@
The validation of inequality (78) does not rely on the any training algorithm and hence it still holds here. To bound the last term in the RHS of (78), we have
$\ \frac{1}{2\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha\nabla f(W_{k};\xi_{k})\|^{2}]$
$=$ $\ \frac{\alpha}{2}\mathbb{E}_{\xi_{k}}[\|\tilde{\nabla}^{k}f-\nabla f(W_{k};\xi_{k})\|^{2}]$
$\leq$ $\ \frac{\alpha C_{f}}{2}\mathbb{E}_{\xi_{k}}\left[\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|\right]^{2}$
$\leq$ $\ \frac{\alpha C_{f}M^{2}}{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}]$ (98)
@@
$\ -\frac{1}{4\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]$ (99)
where the second inequality is earned by plugging in (F).
Construct a Lyapunov function as
@@
$\sum_{m^{\prime}=1}^{M}{\mathbb{E}}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k}\|^{2}]$ $=\sum_{m^{\prime}=1}^{M}{\mathbb{E}}_{\xi_{k}}[\|W_{k+1}^{(m^{\prime})}-W_{k}^{(m^{\prime})}\|^{2}]$
$={\mathbb{E}}_{\xi_{k}}[\|W_{k+1}-W_{k}\|^{2}]$
@@
$\leq 2{\mathbb{E}}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]+2\alpha^{2}\sigma^{2}.$ (101)
and the last inequality is earned by letting $\alpha\leq\sqrt{\frac{1}{4C_{f}M^{3}}}$. Reorganizing the terms yield
@@
$\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}}[\|\nabla f(W_{k})\|^{2}]\leq$ $\ \frac{2({\mathbb{V}}_{0}-{\mathbb{E}}[{\mathbb{V}}_{K}])}{\alpha K}+2\alpha L_{f}\sigma^{2}+2\alpha^{2}C_{f}M^{3}\sigma^{2}$
@@
The proof is completed. ∎
layer name output size 10-layer 34-layer
conv1 112$\times$112 7$\times$7, 64, stride 2
conv2_x 56$\times$56 3$\times$3 max pool, stride 2
$\left[\begin{array}[]{c}\text{3$\times$3, 64}\\[-1.00006pt] \text{3$\times$3, 64}\end{array}\right]$$\times$1 $\left[\begin{array}[]{c}\text{3$\times$3, 64}\\[-1.00006pt] \text{3$\times$3, 64}\end{array}\right]$$\times$3
conv3_x 28$\times$28 $\left[\begin{array}[]{c}\text{3$\times$3, 128}\\[-1.00006pt] \text{3$\times$3, 128}\end{array}\right]$$\times$1 $\left[\begin{array}[]{c}\text{3$\times$3, 128}\\[-1.00006pt] \text{3$\times$3, 128}\end{array}\right]$$\times$4
conv4_x 14$\times$14 $\left[\begin{array}[]{c}\text{3$\times$3, 256}\\[-1.00006pt] \text{3$\times$3, 256}\end{array}\right]$$\times$1 $\left[\begin{array}[]{c}\text{3$\times$3, 256}\\[-1.00006pt] \text{3$\times$3, 256}\end{array}\right]$$\times$6
conv5_x 7$\times$7 $\left[\begin{array}[]{c}\text{3$\times$3, 512}\\[-1.00006pt] \text{3$\times$3, 512}\end{array}\right]$$\times$1 $\left[\begin{array}[]{c}\text{3$\times$3, 512}\\[-1.00006pt] \text{3$\times$3, 512}\end{array}\right]$$\times$3
fc 1$\times$1 average pool, fully connected, softmax
block number 6 18
Table 3: ResNet models in the simulations. Downsampling is performed by conv3_1, conv4_1, and conv5_1 with a stride of 2.
Appendix G Simmulation details
This section provides details about the experiments in Section 5. The analog training algorithm, i.e., Analog SGD, is provided by the open-source simulation toolkit AIHWKIT [31], which has Apache-2.0 license; see github.com/IBM/aihwkit. We use Softbound device provided by AIHWKIT to simulate the asymmetric linear device (ALD), by setting its upper and lower bound as $\tau$. The digital algorithm, including SGD, and datasets used in this paper, including CIFAR10 and CIFAR-100, are provided by PyTorch, which has BSD license; see https://github.com/pytorch/pytorch.
The other setting follows the standard settings of AIHWKIT, including output noise (0.5 % of the quantization bin width), quantization and clipping (output range set 20, output noise 0.1, and input and output quantization to 8 bits). Noise and bound management techniques are used in [33]. A learnable scaling factor is set after each analog layer, which is updated using digital SGD.
Compute Resources. We conduct our simulations on an NVIDIA RTX 3090 GPU, which has 24GB memory. The simulations take from 30 minutes to 7 hours, depending on the size of the model and the dataset.
Statistical Significance. The simulations reported in Figure 2 and 4 are repeated three times. The randomness originates from the data shuffling, random initialization, and random noise in the analog device simulator. The mean and standard deviation are calculated from the final output of each run using NumPy library.
Image classification on CIFAR10/CIFAR100 with Resnet models. In the simulations, we train ResNet models on two benchmark datasets: CIFAR-10 and CIFAR-100. The CIFAR-10 dataset consists of 60,000 32$\times$32 color images across 10 classes, with each class representing a distinct category such as airplanes, cars, and birds. The CIFAR-100 dataset is similar in format but includes 100 classes.
In Figure 4 and 5, ResNet10 is chosen while in Figure 1, a larger model, ResNet34 is chosen. As Table 3 shows, ResNet10 and ResNet34 are composed of 6 and 18 blocks, respectively. In Figure 4 and 5, each stage contains one block while in Figure 1, 18 blocks are separated into $M=1,2,4,8,12,16$ stages as evenly as possible.
Each model was trained over 300 epochs using a mini-batch size of $B_{\text{mini}}=128$ and micro-batch-size $B_{\text{micro}}=16$ by default. The initial learning rate of $\alpha=0.1$ was employed, while for Resnet34 with stage number $M=12,16$, the initial learning rate is $\alpha=0.01$. The transfer learning rate of Tiki-Taka is chosen to be equal to the initial learning rate and kept fit during the training. The learning rate $\alpha$ was reduced by a factor of 10 every 100 epochs. To further enhance the model’s ability to generalize across diverse images, we applied data augmentation techniques such as random cropping, horizontal flipping, and color jittering [34].

### v2

@@ Appendix E Proof of Theorem 2: Analog SGD with asynchronous pipeline
$\ -\left(\frac{1}{4\alpha}-\alpha(1+\frac{1}{u})(1+W_{\max,\infty}/\tau)^{2}C_{f}^{2}M^{2}\right)\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]$
$\ +\alpha^{3}(1+\frac{1}{u})(1+W_{\max,\infty}/\tau)^{2}C_{f}^{2}M^{3}\sigma^{2}.$
Taking expectation, averaging (109) over $k$ from 0 to K1K-1, and choosing the stepsize $\alpha\leq\frac{1}{2\sqrt{1+1/u}(1+W_{\max,\infty}/\tau)C_{f}M}$ yield
@@
$\leq$ $\ \frac{2({\mathbb{V}}_{0}-{\mathbb{E}}[{\mathbb{V}}_{k+1}])}{\alpha K(1-(1+u)W_{\max,\infty}^{2}/\tau^{2})}+\frac{2\alpha L_{f}\sigma^{2}}{1-(1+u)W_{\max,\infty}^{2}/\tau^{2}}$
$\ +\sigma^{2}\frac{1}{K}\sum_{k=0}^{K-1}\frac{(1+u)\|W_{k}\|^{2}_{\infty}/\tau}{1-(1+u)\|W_{k}\|^{2}_{\infty}/\tau^{2}}+\frac{\alpha^{2}(1+1/u)(1+W_{\max,\infty}/\tau)^{4}C_{f}^{2}M^{3}\sigma^{2}}{(1-(1+u)W_{\max,\infty}^{2}/\tau^{2})}.$
@@
$\leq$ $\ 4\sqrt{\frac{(f(W_{0})-f^{*})\sigma^{2}L_{f}}{K}}\frac{1}{1-(1+u)W_{\max,\infty}^{2}/\tau^{2}}+\sigma^{2}S_{K}^{\prime}+{\mathcal{O}}\left(\frac{1+1/u}{K}\right).$
@@
The proof is completed now. ∎
Appendix F Convergence of Digital Asynchronous Pipeline SGD
Theorem 3 (Iteration complexity, asynchronous pipeline).
@@ Appendix F Convergence of Digital Asynchronous Pipeline SGD
$\ \frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}}\left[\|\nabla f(W_{k})\|^{2}\right]\leq\ 4\sqrt{\frac{(f(W_{0})-f^{*})L_{f}\sigma^{2}}{K}}+\frac{2C_{f}M^{3}(f(W_{0})-f^{*})}{KL_{f}}.$ (113)
@@
The validation of inequality (92) does not rely on any training algorithm, and hence it still holds here. To bound the last term in the RHS of (92), we have
$\ \frac{1}{2\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha\nabla f(W_{k};\xi_{k})\|^{2}]=\frac{\alpha}{2}\mathbb{E}_{\xi_{k}}[\|\tilde{\nabla}^{k}f-\nabla f(W_{k};\xi_{k})\|^{2}]$
$\leq$ $\ \frac{\alpha C_{f}}{2}\mathbb{E}_{\xi_{k}}\left[\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|\right]^{2}\leq\frac{\alpha C_{f}M^{2}}{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}]$ (114)
@@
$\ \mathbb{E}_{\xi_{k}}[f(W_{k+1})]$ (115)
$\leq$ $\ f(W_{k})-\frac{\alpha}{2}\|\nabla f(W_{k})\|^{2}+\alpha^{2}L_{f}\sigma^{2}+\frac{1}{2\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha\nabla f(W_{k};\xi_{k})\|^{2}]$
@@
where the second inequality is earned by plugging in (F). Construct a Lyapunov function as
@@
$\sum_{m^{\prime}=1}^{M}{\mathbb{E}}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k}\|^{2}]$ $=\sum_{m^{\prime}=1}^{M}{\mathbb{E}}_{\xi_{k}}[\|W_{k+1}^{(m^{\prime})}-W_{k}^{(m^{\prime})}\|^{2}]={\mathbb{E}}_{\xi_{k}}[\|W_{k+1}-W_{k}\|^{2}]$ (117)
@@
$\leq 2{\mathbb{E}}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]+2\alpha^{2}\sigma^{2}.$
and the last inequality is earned by letting $\alpha\leq\sqrt{\frac{1}{4C_{f}M^{3}}}$. Reorganizing the terms yields
@@
$\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}}[\|\nabla f(W_{k})\|^{2}]\leq$ $\ \frac{2({\mathbb{V}}_{0}-{\mathbb{E}}[{\mathbb{V}}_{K}])}{\alpha K}+2\alpha L_{f}\sigma^{2}+2\alpha^{2}C_{f}M^{3}\sigma^{2}$ (118)